# A Systematic Evaluation of the Potential of Carbon-Aware Execution for Scientific Workflows

### v3

@@ Abstract
Scientific workflows are widely used to automate scientific data analysis and often involve computationally intensive processing of large datasets on compute clusters. As such, their execution tends to be long-running and resource-intensive, resulting in substantial energy consumption and, depending on the energy mix, carbon emissions. Meanwhile, a wealth of carbon-aware computing methods have been proposed, yet little work has focused specifically on scientific workflows, even though they present a substantial opportunity for carbon-aware computing because they are often significantly delay tolerant, efficiently interruptible, highly scalable and widely heterogeneous.
In this study, we first exemplify the problem of carbon emissions associated with running scientific workflows, and then show the potential for carbon-aware workflow execution. For this, we estimate the carbon footprint of seven real-world Nextflow workflows executed on different cluster infrastructures using both average and marginal carbon intensity data. Furthermore, we systematically evaluate the impact of carbon-aware temporal shifting, and the pausing and resuming of the workflow. Moreover, we apply resource scaling to workflows and workflow tasks. Finally, we report the potential reduction in overall carbon emissions, with temporal shifting capable of decreasing emissions by over 80%, and resource scaling capable of decreasing emissions by 67%.
@@
scientific workflows , carbon-aware computing , carbon footprint , temporal shifting , resource scaling , sustainable computing1 Introduction
@@ 5.2 Sensitivity Analysis
As explained in Section 3, we assumed access to CI forecasts with no error and perfect knowledge of workflow task executions. In practice, CI can be forecasted with a 3–20% error [40], and workflow task runtimes can be predicted with an error of 3–30% [41, 42]. We conducted sensitivity analyses modeling these realistic forecasting and prediction errors considering our entire workflow shifting experiment. Both analyses used average CI data for Great Britain.
@@
We used CI data with a forecast error of 5, 10, and 15%. Fig. 8 shows the results. When the CI forecast had an error, we expected a 2% greater reduction in emissions compared to the optimal reduction using actual data in the 24h window, and a 4.3% greater reduction in the 96h window. While the optimization is sensitive to forecasting errors, the reduction possible remains significant, saving 34–43.5% on average.
@@
We used workflow traces where task runtimes had a prediction error of 10, 15%, and 25%. The results are shown in Table 4. We observed that there was minimal change compared to the optimal reduction in emissions (34% and 43.5%) when applying a workflow task runtime prediction error.
@@
| Runtime Error (%) | 10 | 15 | 25 | 10 | 15 | 25 |
| Reduction Error | 0.03 | 0.04 | 0.03 | 0.00 | 0.01 | 0.01 |

### v4

@@ Abstract
Scientific workflows are critical to scientific data analysis and often involve computationally intensive processing of large datasets on compute clusters. As such, their execution tends to be long-running and resource-intensive, resulting in significant energy consumption and carbon emissions. While carbon-aware computing methods have received considerable attention in general cloud contexts, their application to scientific data analysis workflows remains a critical research gap. Our study addresses this oversight by showing how the delay tolerance, interruptibility, and scalability of scientific workflows can be leveraged for a significantly more sustainable execution model.
In this study, we first quantify the problem of carbon emissions associated with running scientific workflows, and then demonstrate the transformative potential for carbon-aware workflow execution. We estimate the carbon footprint of seven real-world Nextflow workflows executed on diverse dedicated cluster and public cloud resources using high-resolution average and marginal grid carbon intensity data from open and commercial data providers. Furthermore, we conduct a systematic evaluation of the impact of carbon-aware temporal shifting, and the dynamic pausing and resuming of the workflow. Moreover, we investigate the impact of resource scaling at both workflow and workflow task levels. Finally, we report substantial potential reductions in overall carbon emissions, with temporal shifting capable of decreasing emissions by over 80%, and resource scaling by 67%.
@@
scientific workflows , carbon-aware computing , temporal shifting , resource scaling , sustainable computing1 Introduction
@@ 5.2 Sensitivity Analysis
As explained in Section 3, we assumed access to CI forecasts with no error and perfect knowledge of workflow task executions. In practice, CI can be forecasted with a 3–20% error [40], and workflow task runtimes can be predicted with an error of 3–30% [41, 42]. We conducted sensitivity analyses modeling these realistic forecasting and prediction errors considering our entire workflow shifting experiment. Both analyses used average CI data for Great Britain. We report the emissions reduction deviation, which is the shortfall between the emissions reduction achievable with perfect data and the reduction possible when scheduling decisions rely on CI forecasts or task runtime predictions with an error.
@@
We used CI data with a forecast error of 5, 10, and 15%. Table 4 shows the results. When the CI forecast had an error, we saw a 2% reduction deviation in emissions compared to the optimal reduction using actual data in the 24h window, and a 4.3% reduction deviation in the 96h window. While the optimization is sensitive to forecasting errors, the reduction possible remains significant.
| 24h | 96h | |||||
| CI Forecast Error (%) | 5 | 10 | 15 | 5 | 10 | 15 |
| Emissions Reduction Deviation | 0.3 | 0.9 | 2.0 | 0.5 | 1.9 | 4.3 |
@@
We used workflow traces where task runtimes had a prediction error of 10, 15%, and 25%. The results are shown in Table 5. We observed that there was minimal change compared to the optimal reduction in emissions (34% and 43.5%) as a result of workflow task runtime prediction errors.
@@
| Runtime Prediction Error (%) | 10 | 15 | 25 | 10 | 15 | 25 |
| Emissions Reduction Deviation | 0.03 | 0.04 | 0.03 | 0.00 | 0.01 | 0.01 |